### v2

@@ 4 Identification with Strategic Reports
This section studies identification of causal effects when students are strategic in reporting their rankings of acceptable schools. Strategic reports make $P_{j}$ generally different from $Q_{j}$, and $Q_{j}$ is not observed. Identification strategies that condition on the observed $P_{j}$ are potentially invalid for two reasons. First, in contrast to Proposition 1, controlling for $P_{j}$ does not identify the parameters of interest in (1), which condition on $Q_{j}$. Second, controlling for $P_{j}$ breaks the internal validity of the RD in some cases because $P_{j}$ is a variable subject to manipulation by agents.
We propose a two-step identification approach that is robust to strategic reporting. In the first step, the researcher characterizes the set of true local preferences $Q_{j}$ that is compatible with the data and appropriate behavioral assumptions. In the second step, the researcher controls for the constructed local preference sets and partially identifies the parameters in (1). We discuss the first and second steps in Sections 4.1 and 4.2, respectively. Note that our two-step approach differs from the usual two-step control function approach in econometrics. The usual approach is to point-identify the control variable in the first step, while our approach involves partially identifying the control variable.
Section 4.1 presents several tools for the identification of local preference sets. These tools rely on assumptions known to be appropriate in SD and DA contexts, although we do not rule out their applicability in contexts with other mechanisms; e.g., the TTC mechanism satisfies one of the necessary assumptions, such that some of the tools from Section 4.1 are still useful for causal identification of assignment effects under this mechanism. More generally, researchers may utilize preference identification tools that work under alternative assumptions, for example, the methods of agarwal2018 and fack2019. Either way, the researcher must construct a set of local preferences for each individual in the first step.
Section 4.2 describes the second step of our procedure. This step of the procedure features high-level assumptions imposed on the local preference sets such that the researcher is not restricted to the methods in Section 4.1.
@@ 4.1 Partial Identification of Local Preferences
This section provides tools for local preference set identification using assumptions on agents’ behavior and the mechanism. These assumptions are specific to this subsection, and we motivate them with reference to the context of the constrained DA mechanism studied by haeringer2009. haeringer2009 study a game where students submit constrained preference rankings and a mechanism matches students to schools as a function of $P$, $\bm{\mathrm{R}}$, and schools’ capacities. Although the unconstrained DA mechanism is strategy proof, many real-world implementations of DA restrict the number of schools that students can submit in their rankings. In this case, there is a cap $K<J$ such that $1\leq|P|\leq K$, and the submitted ranking $P$ is generally different from the list of acceptable schools in $Q$. When implemented in this way, the DA mechanism is not strategy proof, and there are no dominant strategies. Strategyproofness also breaks down if, instead of facing a cap, students incur an application cost as a function of the number of schools submitted (fack2019).
Lemma 4.2 by haeringer2008 shows that, if a mechanism is strategy proof when $K=J$, then, in the game with $K<J$, any constrained ranking of schools is weakly dominated by the same set of schools ranked according to true preferences. This result implies that a student cannot lose and may possibly gain by taking any arbitrary list with less than or equal to $K$ schools, dropping the unacceptable schools, and ranking the acceptable schools according to her true preferences. A further implication is that if a student’s number of acceptable schools is less than or equal to $K$, then her dominant strategy is to list acceptable schools that reflect her true preferences (Proposition 4.2 of haeringer2009). These implications give rise to a class of undominated strategies according to the following definitions of partial order.
@@
Lemma 2 in Section A.3 of the appendix summarizes the implications of the result on partial orders from haeringer2008; haeringer2009 in terms of our Definition 6. In short, for a student with true preferences $Q$, any $P$ is weakly dominated by a weak partial order $P^{*}$ of $Q$ that has the same acceptable schools as $P$; in turn, $P^{*}$ is weakly dominated by a strong partial order $P^{**}$ of $Q$ that contains the same set of acceptable schools as $P^{*}$. Every strong partial order is a weak partial order, but the converse is not true. Our definition of a weak partial order strategy is similar to the definition of the dropping strategy in kojima2009incentives.
Assuming that agents always submit a strong partial order implies they reveal their true ordered list of acceptable schools whenever they submit $P$ with fewer schools than the cap $K$. This could be a strong behavioral assumption in some contexts where agents have more than $K$ acceptable schools but have a strong preference for admission to a smaller-than-$K$ set of schools. In this case, they may submit $|P|<K$ not because it reflects their full list of acceptable schools but simply because they may not want to incur the mental costs of ranking all schools up to $K$. In the rest of this subsection, we consider mechanisms that impose a cap $K$ on $P$ and assume that students submit a weak partial order of their true preferences.
@@
An assumption maintained throughout this paper is that the cutoff characterization from Definition 2 applies to the mechanism. This assumption says that $\mu(\omega)=Q(B(\bm{\mathrm{S}}(\omega))$ for every $\omega\in\Omega$. azevedo2016 show that stability is equivalent to cutoff characterization with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and cutoffs that equal the minimum score of the admitted students in each school. Thus, it is worth discussing the stability of the constrained DA mechanism. Theorem 6.3 in haeringer2009 demonstrates that any Nash equilibrium in constrained DA where $\bm{\mathrm{R}}$ satisfies Ergin acyclicity leads to a stable matching in the finite economy.888 Ergin acyclicity ensures that no student can block a potential improvement for any two other students without affecting her own assignment. See ergin2002 for the formal definition. Even without Ergin acyclicity, some Nash equilibria still produce stability. SD always satisfies Ergin acyclicity, and thus every Nash equilibrium in SD produces a stable matching. fack2019 also study the constrained DA mechanism. They extend Theorem 6.3 from haeringer2009 to (pure-strategy) Bayesian Nash equilibria in the continuum economy (Proposition A3, Online Appendix A.2.5, fack2019). fack2019 also provide primitive conditions for finite economies where students play partial orders to converge to a continuum economy with a stable equilibrium (Proposition 5, fack2019). They further provide a test for implications of stability and find no empirical or simulation evidence against it. In the context of the constrained TTC mechanism, any Nash equilibrium leads to a stable matching as long as $\bm{\mathrm{R}}$ satisfies Kesten acyclicity (Theorem 6.4 in haeringer2009). Therefore, constrained SD, DA, and TTC all satisfy the assumption of cutoff characterization as in Definition 2 with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ under the appropriate conditions.
@@
The DA mechanism, constrained or unconstrained, produces a matching that is stable wrt reported preferences $P$. Stability wrt $P$ leads to a cutoff characterization wrt $P$ according to the work of azevedo2016. This cutoff characterization has scores $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and admission cutoffs that equal the smallest scores of admitted students in each school. In other words, this is the same cutoff characterization from Definition 2 except that $Q$ is replaced with $P$. Cutoff characterization wrt $P$ is natural in DA but not necessarily in other mechanisms, so we state it in the following assumption.
@@
Assumption 4 essentially says that agents are matched to their best feasible options, where best is now defined according to $P$. Assumption 4 is convenient because it allows us to write a simple expression for the identified set of local preferences in Proposition 2 below; However, it is not a necessary assumption for the identification of those sets. The convenience comes from the fact that Assumption 4 implies $\mu=P(B(\bm{\mathrm{S}}))=Q(B(\bm{\mathrm{S}}))$, where both $\mu$ and $P$ are observed and $P$ and $Q$ are related via the weak partial order assumption. If we drop Assumption 4, we have only one equality $\mu=Q(B(\bm{\mathrm{S}}))$, which leads to larger sets of local preferences in Proposition 2 below. This is useful to know for settings such as those with the TTC mechanism, which is not stable wrt $P$ and thus is not a natural candidate for study with an identification strategy relying on Assumption 4.
@@
Suppose Assumptions 3 and 4 hold. Select a school $j$ with cutoff $c_{j}$. Consider a student with scores $\bm{\mathrm{S}}\equiv(S_{j},\bm{\mathrm{S}}_{-j})$ and submitted preferences $P$. Call $(a,b)=P_{j}$. For this student, define as $N_{j}^{+}=B^{+}_{j}(\bm{\mathrm{S}})\setminus\left\{P\cup\{0\}\right\}$ and $N_{j}^{-}=B^{-}_{j}(\bm{\mathrm{S}})\setminus\left\{P\cup\{0\}\right\}$, respectively, the sets of unlisted feasible schools in the counterfactual budget sets to the right and the left of the cutoff. Then, the $Q_{j}$ of this student belongs to $\bm{\mathrm{Q}}_{j}$, where the set $\bm{\mathrm{Q}}_{j}$ is defined as follows:
$\bm{\mathrm{Q}}_{j}=\left\{\begin{array}[]{ll}\{(a,b)\},&\text{ if }S_{j}\geq c_{j}\text{ and }a=b,\\ \{(a,b)\}\cup\left(\{a\}\times N_{j}^{-}\right),&\text{ if }S_{j}\geq c_{j}\text{ and }a\neq b,\\ \{(a,b)\}\cup\left((N_{j}^{+}\setminus N_{j}^{-})\times\{b\}\right)&\text{ if }S_{j}<c_{j},\end{array}\right.$ (5)
where ($\left(\{a\}\times N_{j}^{-}\right)$ denotes the set formed by the Cartesian product of $a$ and elements in $N_{j}^{-}$ and ($\left(\{a\}\times N_{j}^{-}\right)=\emptyset$ if $N_{j}^{-}=\emptyset$.
@@
$\bm{\mathrm{Q}}_{j}=\left\{\begin{array}[]{ll}\{(a,b)\},&\text{ if }|P|<K,\text{ or if }|P|=K,~{}S_{j}\geq c_{j},\text{ and }a=b,\\ \{(a,b)\}\cup\left(\{a\}\times N_{j}^{-}\right),&\text{ if }|P|=K,~{}S_{j}\geq c_{j},\text{ and }a\neq b,\\ \{(a,b)\}\cup\left((N_{j}^{+}\setminus N_{j}^{-})\times\{b\}\right)&\text{ if }|P|=K\text{ and }S_{j}<c_{j}.\end{array}\right.$ (9)
Finally, the characterization in (5) is sharp if the distribution of $Q$ conditional on $P$ and $\bm{\mathrm{S}}$ has full support, that is, if every $Q$ that satisfies Assumptions 3 and 4 is in that support. Likewise, (9) is sharp if the distribution of $Q$ conditional on $P$ and $\bm{\mathrm{S}}$ has full support under Assumptions 3 and 4 and $P$ being a strong partial order.
Example (SD Example, Continued).
Suppose the cap constraint is $K=3$ and the four schools are acceptable for everyone. We consider all agents whose $P_{4}=(4,2)$. For example, if agents submit strong partial orders, they submit either $P=\{4,2,1\}$ or $P=\{4,2,3\}$. The assumption of cutoff characterization wrt $P$ (Assumption 4) says that these agents are matched to school $4$ if $S_{1}\geq c_{4}$ and to school $2$ otherwise. To keep things simple, consider five different types of true preferences: $Q^{(1)}=\{4,2,3,1,0\}$, $Q^{(2)}=\{4,3,2,1,0\}$, $Q^{(3)}=\{4,1,3,2,0\}$, $Q^{(4)}=\{3,4,2,1,0\}$, and $Q^{(5)}=\{1,3,2,4,0\}$. The weak partial order assumption rules out $Q^{(5)}$ because $42Q^{(5)}$ 4 contradicts $4$ being reported preferred to $2$. The maintained assumption of cutoff characterization (Definition 2) further rules out more types of $Q$, depending on whether $S_{1}\geq c_{4}$ or $S_{1}<c_{4}$:
@@
This example illustrates why an RD that controls only for $P_{4}=(4,2)$ may fail to identify the treatment effect of a change in assignment from school $4$ to school $2$. As the score $S_{1}$ crosses the cutoff $c_{4}$, the support of the distribution of true local preference types generally changes discontinuously from having only $(4,2)$ to having $(4,1)$, $(4,2)$, and $(4,3)$. This possibility invalidates RD identification because true preferences and potential outcomes are plausibly dependent on each other; in this case, the distribution of potential outcomes conditional on $S_{1}=s$ and $P_{4}=(4,2)$ changes discontinuously at $S_{1}=c_{4}$. For the possibility to arise, a positive mass of agents must change their submission $P$ as their score $S_{1}$ crosses $c_{4}$. This does not necessarily require that these agents are certain ex ante about the cutoff value $c_{4}$ but simply that they have expectations that change as $S_{1}$ crosses $c_{4}$.
Proposition 2 identifies all possible values of $Q_{j}$ for students near a cutoff $c_{j}$ as a function of their scores and submitted preferences. In some settings, many unlisted programs may be feasible for students, which translates into sets $\bm{\mathrm{Q}}_{j}$ with many values. For example, the Chilean data have $K=8$ but more than 1,000 options; a student may have numerous feasible options but not consider many of them. It is possible to obtain smaller sets $\bm{\mathrm{Q}}_{j}$ by imposing further assumptions on the expectations that students have when they submit $P$.
@@
agarwal2018 propose a general framework to rationalize strategic reporting as the optimal solution to an expected utility maximization problem. In this framework, agents have private information about their preferences and scores and form beliefs about the distribution of other people’s preferences and scores. These beliefs plus knowledge of the mechanism lead the rational agent to derive probabilities of admission to the various schools as a function of the agent’s private information and expectations about other agents. The agent then chooses the submission $P$ that maximizes her expected utility.
@@
We assume that agents are expected utility maximizers and note that it is sufficient for each individual to consider beliefs on admission cutoffs. This is the case in our continuum economy with cutoff characterization because cutoffs and scores fully characterize the agent’s budget set. Given the student’s scores, a distribution of possible cutoffs translates into a distribution of possible budget sets. Under Assumption 4, the student is admitted to the best school according to the submission $P$ among the available schools in the budget set. Therefore, beliefs on cutoffs translate into probabilities of admission to various schools for any given $P$. We make an assumption on the distribution of cutoffs expected by agents that has to do with the concept of uniformly more accessible schools.
@@
$\{s_{d}\geq C_{d}\}\subseteq\{s_{e}\geq C_{e}\}.$
Second, for any two fixed (i.e., nonrandom) submissions $P$ and $\widetilde{P}$ such that $P$ has $d$ but does not have $e$ and $\widetilde{P}$ equals $P$ except for $e$ in the place of $d$, there exists $u\in\{0,1,\ldots,|P|\}$ for which
@@
Suppose Assumptions 35 hold. Consider a student with reported preference ranking $P$. Let $\left(P\times P^{c}\right)$ be the Cartesian product of listed and unlisted schools, respectively, $P$ and $P^{c}$. If $(d,e)\in UMAS\cap\left(P\times P^{c}\right)$, then $dQe$.
@@
$\bm{\mathrm{Q}}_{j}=\left\{\begin{array}[]{ll}\{(a,b)\},&\text{ if }S_{j}\geq c_{j}\text{ and }a=b,\\ \{(a,b)\}\cup\left(\{a\}\times A_{j}^{-}\right),&\text{ if }S_{j}\geq c_{j}\text{ and }a\neq b,\\ \{(a,b)\}\cup\left(A_{j}^{+}\times\{b\}\right)&\text{ if }S_{j}<c_{j}.\end{array}\right.$ (13)
@@
$\bm{\mathrm{Q}}_{j}=\left\{\begin{array}[]{ll}\{(a,b)\},&\text{ if }|P|<K,\text{ or if }|P|=K,~{}S_{j}\geq c_{j},\text{ and }a=b,\\ \{(a,b)\}\cup\left(\{a\}\times A_{j}^{-}\right),&\text{ if }|P|=K,~{}S_{j}\geq c_{j},\text{ and }a\neq b,\\ \{(a,b)\}\cup\left(A_{j}^{+}\times\{b\}\right)&\text{ if }|P|=K\text{ and }S_{j}<c_{j}.\end{array}\right.$ (17)
@@
Example (SD Example, Continued).
The set of uniformly more accessible schools is $UMAS=\{(2,1),(3,2),(3,1),(4,3),(4,2),(4,1)\}$. Assumption 5 shrinks the set $\bm{\mathrm{Q}}_{4}$ of those agents with $S_{1}\geq c_{4}$ and $P=\{4,2,3\}$. Applying the assumption changes $\bm{\mathrm{Q}}_{4}=\{(4,2),(4,1)\}$ to $\bm{\mathrm{Q}}_{4}=\{(4,2)\}$ because $1$ is uniformly more accessible than $2$, $2$ is listed, and $1$ is not listed, so Proposition 3 implies $12Q$ 1.

### v3

@@ 4 Identification with Strategic Reports
This section studies identification of causal effects when students are strategic in reporting their rankings of acceptable schools. Strategic reports make $P_{j}$ generally different from $Q_{j}$, and $Q_{j}$ is not observed. In contrast to Proposition 1, controlling for $P_{j}$ does not identify the parameters of interest in (3), which condition on $Q_{j}$. Moreover, controlling for $P_{j}$ may break the internal validity of the RD in some cases because $P_{j}$ is a variable subject to manipulation by agents. The second reason is more subtle and we explain it in Section 4.2 below in terms of our recurring SD Example (Part III).
We propose a two-step identification approach that is robust to strategic reporting. In the first step, the researcher characterizes the set of true local preferences $Q_{j}$ that is compatible with the data and appropriate behavioral assumptions. In the second step, the researcher controls for the constructed local preference sets and partially identifies the parameters in (3). We discuss the first and second steps in Sections 4.1 and 4.2, respectively. Note that our two-step approach differs from the usual two-step control function approach in econometrics. The usual approach is to point-identify the control variable in the first step, while our approach involves partially identifying the control variable. Thus, we refer to our two-step procedure as a control mapping approach.
Section 4.1 presents several tools for the identification of local preference sets. These tools rely on assumptions known to be appropriate in SD and DA contexts, although we do not rule out their applicability in contexts with other mechanisms; e.g., the TTC mechanism satisfies one of our assumptions, such that some of the tools from Section 4.1 are still useful. More generally, researchers may utilize preference identification tools that work under alternative assumptions, for example, the methods of Agarwal and Somaini (2018) and Fack et al. (2019). Either way, the researcher must construct a set of local preferences for each individual in the first step.
Section 4.2 describes the second step of our procedure. This step features high-level assumptions imposed on the local preference sets such that the researcher is not restricted to the methods of Section 4.1.
@@ 4.1 Partial Identification of Local Preferences
This section provides tools for set identification of local preferences using assumptions on agents’ behavior and the mechanism. These assumptions are specific to this subsection, and we motivate them with reference to the context of the constrained DA mechanism studied by Haeringer and Klijn (2009). Haeringer and Klijn (2009) study a game where students submit constrained preference rankings and a mechanism matches students to schools as a function of $P$, $\bm{\mathrm{R}}$, and schools’ capacities. Although the unconstrained DA mechanism is strategy proof, many real-world implementations of DA restrict the number of schools that students can submit in their rankings. In this case, there is a cap $K<J$ such that $1\leq|P|\leq K$, and the submitted ranking $P$ is generally different from the list of acceptable schools in $Q$. When implemented in this way, the DA mechanism is not strategy-proof, and there are no dominant strategies. Strategyproofness also breaks down if, instead of facing a cap, students incur an application cost as a function of the number of schools submitted (Fack et al., 2019).
Lemma 4.2 by Haeringer and Klijn (2008) shows that, if a mechanism is strategy proof when $K=J$, then, in the game with $K<J$, any constrained ranking of schools is weakly dominated by the same set of schools ranked according to true preferences. This result implies that a student cannot lose and may possibly gain by taking any arbitrary list with less than or equal to $K$ schools, dropping the unacceptable schools, and ranking the acceptable schools according to her true preferences. A further implication is that if a student’s number of acceptable schools is less than or equal to $K$, then her dominant strategy is to submit her true list of acceptable schools (Proposition 4.2 of Haeringer and Klijn (2009)). These implications give rise to a class of undominated strategies according to the following definitions of partial order.
@@
Lemma 2 in Section A.2 of the appendix summarizes the implications of the result on partial orders from Haeringer and Klijn (2008, 2009) in terms of our Definition 6. In short, for a student with true preferences $Q$, any $P$ is weakly dominated by a weak partial order $P^{*}$ of $Q$ that has the same acceptable schools as $P$; in turn, $P^{*}$ is weakly dominated by a strong partial order $P^{**}$ of $Q$ that contains the same set of acceptable schools as $P^{*}$. Every strong partial order is a weak partial order, but the converse is not true. Our definition of a weak partial order strategy is similar to the definition of the dropping strategy from Kojima and Pathak (2009).
Assuming that agents always submit a strong partial order implies they reveal their true ordered list of acceptable schools whenever they submit $P$ with fewer schools than the cap $K$. This could be a strong behavioral assumption in some contexts where agents have more than $K$ acceptable schools but have a strong expectation that they will gain admission to a smaller-than-$K$ set of schools. In this case, they may submit $|P|<K$ not because it reflects their full list of acceptable schools but simply because they may not want to incur the all the costs of ranking all schools up to $K$. In the rest of this subsection, we consider mechanisms that impose a cap $K$ on $P$ and assume that students submit a weak partial order of their true preferences.
@@
An assumption maintained throughout this paper is that the cutoff characterization from Definition 2 applies to the mechanism. This assumption says that $\mu(\omega)=Q(B(\bm{\mathrm{S}}(\omega))$ for every $\omega\in\Omega$. Azevedo and Leshno (2016) show that stability is equivalent to cutoff characterization with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and cutoffs that equal the minimum score of the admitted students in each school. Thus, it is worth discussing the stability of the constrained DA mechanism. Theorem 6.3 from Haeringer and Klijn (2009) demonstrates that any Nash equilibrium in constrained DA where $\bm{\mathrm{R}}$ satisfies Ergin acyclicity leads to a stable matching in the finite economy.999 Ergin acyclicity ensures that no student can block a potential improvement for any two other students without affecting her own assignment. See Ergin (2002) for the formal definition. Even without Ergin acyclicity, some Nash equilibria still produce stability. SD always satisfies Ergin acyclicity, and thus every Nash equilibrium in SD produces a stable matching. Fack et al. (2019) also study the constrained DA mechanism. They extend Theorem 6.3 from Haeringer and Klijn (2009) to (pure-strategy) Bayesian Nash equilibria in the continuum economy (Proposition A3, Online Appendix A.2.5, Fack et al. (2019)). Fack et al. (2019) also provide primitive conditions for finite economies where students play partial orders to converge to a continuum economy with a stable equilibrium (Proposition 5, Fack et al. (2019)). They further provide a test for implications of stability and find no empirical or simulation evidence against it. In the context of the constrained TTC mechanism, any Nash equilibrium leads to a stable matching as long as $\bm{\mathrm{R}}$ satisfies Kesten acyclicity (Theorem 6.4 from Haeringer and Klijn (2009)). Therefore, constrained SD, DA, and TTC all satisfy the assumption of cutoff characterization as in Definition 2 with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ under the appropriate conditions.
@@
The DA mechanism, constrained or unconstrained, produces a matching that is stable wrt reported preferences $P$. Stability wrt $P$ leads to a cutoff characterization wrt $P$ according to the work of Azevedo and Leshno (2016). This cutoff characterization has scores $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and admission cutoffs that equal the smallest scores of admitted students in each school. In other words, this is the same cutoff characterization from Definition 2 except that $Q$ is replaced with $P$. Cutoff characterization wrt $P$ is natural in DA but not necessarily in other mechanisms, so we state it in the following assumption.
@@
Assumption 4 essentially says that agents are matched to their best feasible options, where best is now defined according to $P$. Assumption 4 is convenient because it allows us to write a simple expression for the identified set of local preferences in Proposition 2 below; however, it is not a necessary assumption for the identification of those sets. The convenience comes from the fact that Assumption 4 implies $\mu=P(B(\bm{\mathrm{S}}))=Q(B(\bm{\mathrm{S}}))$, where both $\mu$ and $P$ are observed and $P$ and $Q$ are related via the weak partial order assumption. If we drop Assumption 4, we have only one equality $\mu=Q(B(\bm{\mathrm{S}}))$, which leads to larger sets of local preferences in Proposition 2 below. This is useful to know for settings such as those with the TTC mechanism, which is not stable wrt $P$.
@@
Suppose Assumptions 3 and 4 hold. Select a school $j$ with cutoff $c_{j}$. Consider a student with scores $\bm{\mathrm{S}}\equiv(S_{j},\bm{\mathrm{S}}_{-j})$ and submitted preferences $P$. Call $(a,b)=P_{j}$. For this student, define $N_{j}^{+}=B^{+}_{j}(\bm{\mathrm{S}})\setminus\left\{P\cup\{0\}\right\}$ and $N_{j}^{-}=B^{-}_{j}(\bm{\mathrm{S}})\setminus\left\{P\cup\{0\}\right\}$, respectively, the sets of unlisted feasible schools in the counterfactual budget sets to the right and the left of the cutoff. Then, the $Q_{j}$ of this student belongs to $\bm{\mathrm{Q}}_{j}$, where the set $\bm{\mathrm{Q}}_{j}$ is defined as follows:
[…]
where $\left(\{a\}\times N_{j}^{-}\right)$ denotes the set formed by the Cartesian product of $a$ and elements in $N_{j}^{-}$ and $\left(\{a\}\times N_{j}^{-}\right)=\emptyset$ if $N_{j}^{-}=\emptyset$.
@@
[…]
Finally, the characterization in (7) is sharp if the distribution of $Q$ conditional on $P$ and $\bm{\mathrm{S}}$ has full support, that is, if every $Q\in\mathcal{Q}$ that satisfies Assumptions 3 and 4 is in that support. Likewise, (11) is sharp if the distribution of $Q$ conditional on $P$ and $\bm{\mathrm{S}}$ has full support under Assumptions 3 and 4 and $P$ being a strong partial order.
We illustrate the proposition in terms of the SD Example.
Example (SD Example, Part III).
Suppose the cap constraint is $K=3$ and the four schools are acceptable for everyone. We consider all agents whose $P_{4}=(4,2)$. For example, if agents submit strong partial orders, they submit either $P=\{4,2,1\}$ or $P=\{4,2,3\}$. The assumption of cutoff characterization wrt $P$ (Assumption 4) says that these agents are matched to school $4$ if $S_{1}\geq c_{4}$ and to school $2$ otherwise. To keep things simple, consider five different types of true preferences: $Q^{(1)}=\{4,2,3,1,0\}$, $Q^{(2)}=\{4,3,2,1,0\}$, $Q^{(3)}=\{4,1,3,2,0\}$, $Q^{(4)}=\{3,4,2,1,0\}$, and $Q^{(5)}=\{1,3,2,4,0\}$. The weak partial order assumption rules out $Q^{(5)}$ because $2Q^{(5)}4$ contradicts $4$ being reported preferred to $2$. The maintained assumption of cutoff characterization (Definition 2) further rules out more types of $Q$, depending on whether $S_{1}\geq c_{4}$ or $S_{1}<c_{4}$:
@@
This example illustrates why an RD at $c_{4}$ that controls for $P_{4}=(4,2)$ might be problematic. The range of possibilities for true preference types is different between individuals above and below the cutoff. We see types $(4,1)$, $(4,2)$, and $(4,3)$ above the cutoff but only type $(4,2)$ below the cutoff. Suppose in an extreme case there is a nonzero fraction of individuals with types $(4,1)$ or $(4,3)$ right above the cutoff. True preferences arguably affect outcomes and this implies that the distribution of outcomes conditional on $P_{4}=(4,2)$ and $S_{1}=s$ changes as $s$ crosses the cutoff $c_{4}$ even without a treatment effect. This extreme case breaks internal validity of RD identification. Even if the fraction of types $(4,1)$ or $(4,3)$ at the cutoff is zero and RD identification is valid, a continuous but sudden increase in the fraction of such agents on the right of the cutoff may bring severe bias issues and make inference impossible (e.g., Bertanha and Moreira (2020), in particular, Sections 5 and A.6).
For the issue to arise when we control for $P_{4}=(4,2)$, there must be at least a fraction of agents with true local preferences $(4,1)$ and $(4,3)$ who change their submission behavior discontinuously as a function of $S_{1}$ as $S_{1}$ crosses the value of $c_{4}$; i.e., they must change $P$ such that $P_{4}\neq(4,2)$ below the cutoff and $P_{4}=(4,2)$ above the cutoff. Intuitively, this requires these agents to have good ex-ante knowledge about the ex-post value of the cutoff $c_{4}$. The more knowledge they have about the ex-post cutoff values, the sharper will be their change in behavior around the cutoff, and the closer we are to the extreme case mentioned above. Section A.5 in the appendix presents a numerical example of an economy with agents that maximize expected utility in face of cutoff uncertainty and exhibit such discontinuous behavior at the ex-post cutoffs. Finally, Section 5.1 displays empirical evidence from Chile that supports agents having good a priori knowledge of cutoffs to the point of influencing their submissions.
Proposition 2 identifies all possible values of $Q_{j}$ for students near a cutoff $c_{j}$ as a function of their scores and submitted preferences. In some contexts, students may have a large number of feasible but unlisted programs, resulting in sets $\bm{\mathrm{Q}}_{j}$ with many possible values. For instance, in the Chilean data, $K=8$ but there are over 1,000 programs; a student may have many feasible options but choose not to list most of them. We now introduce one approach to reducing the size of $\bm{\mathrm{Q}}_{j}$ by placing additional restrictions on the expectations students hold when submitting $P$. Other context-specific assumptions can also be employed—for example, in Section 5, we use an assumption based on preference for fields of study when analyzing post-secondary education in Chile.
@@
Agarwal and Somaini (2018) propose a general framework to rationalize strategic reporting as the optimal solution to an expected utility maximization problem. In this framework, agents have private information about their preferences and scores and form beliefs about the distribution of other people’s preferences and scores. These beliefs plus knowledge of the mechanism lead the rational agent to derive probabilities of admission to the various schools as a function of the agent’s private information and expectations about other agents. The agent then chooses the submission $P$ that maximizes her expected utility.
@@
For our next proposition, we assume that agents are expected utility maximizers where the uncertainty about their match comes from uncertainty about what the admission cutoffs will be after the matching algorithm is run. As such, we assume each agent forms beliefs on admission cutoffs (see Section A.4 in the appendix for a formal definition of the problem of the agent). Uncertainty about cutoffs is key in our continuum economy with cutoff characterization because cutoffs and scores fully characterize the agent’s budget set. Given the student’s scores, a distribution of possible cutoffs translates into a distribution of possible budget sets. Under Assumption 4, the student is admitted to the best school according to the submission $P$ among the available schools in the budget set. Therefore, beliefs on cutoffs translate into probabilities of admission to various schools for any given $P$. We make an assumption on the distribution of cutoffs expected by agents that has to do with the concept of uniformly more accessible schools.
@@
$\{s_{d}\geq C_{d}\}\subseteq\{s_{e}\geq C_{e}\},$
and second, for any two fixed (i.e., nonrandom) submissions $P$ and $\widetilde{P}$ such that $P$ has $d$ but does not have $e$ and $\widetilde{P}$ equals $P$ except for $e$ in the place of $d$, there exists $u\in\{0,1,\ldots,|P|\}$ for which
@@
Suppose Assumptions 35 hold. Consider a student with reported preference ranking $P$. Let $\left(P\times P^{c}\right)$ be the Cartesian product of listed and unlisted schools, respectively, $P$ and $P^{c}$. If $(d,e)\in UMAS\cap\left(P\times P^{c}\right)$, then dQedQe.
@@
[…]
@@
[…]
@@
Example (SD Example, Part IV).
The set of uniformly more accessible schools is $UMAS=\{(2,1),(3,2),(3,1),(4,3),(4,2),(4,1)\}$. Assumption 5 shrinks the set $\bm{\mathrm{Q}}_{4}$ of those agents with $S_{1}\geq c_{4}$ and $P=\{4,2,3\}$. Applying the assumption changes $\bm{\mathrm{Q}}_{4}=\{(4,2),(4,1)\}$ to $\bm{\mathrm{Q}}_{4}=\{(4,2)\}$ because $1$ is uniformly more accessible than $2$, 22 is listed, and $1$ is not listed, so Proposition 3 implies $2Q1$.